# Stochastic LQR Design With Disturbance Preview

### v4

@@ Abstract
Index Terms:
Stochastic LQR, discrete time, disturbance preview
@@ I Introduction
This paper considers control design for a discrete-time, stochastic Linear Quadratic Regulator (LQR) problem with finite preview information of the disturbance. We consider both the finite and infinite horizon problems. Our specific formulation is motivated by recent work on regret-based control [1, 2, 3, 4, 5, 6, 7]. These regret-based formulations often use a controller with full knowledge of the disturbance as a baseline for performance. The optimal controller with finite preview, as derived in our paper, can provide an alternative baseline for regret-based formulations.
@@
Two of the most related works are [14] and [24]. Calculus of variations is used in [14] to derive a deterministic preview control solution. The approach starts by deriving the optimal non-causal controller with full knowledge of future disturbance, similar to that given in Theorem 11.2.1 of [29]. This optimal non-causal controller is then converted to a finite-preview controller by taking the expectation for unavailable future (post-preview) information which is zero. In [24], an LQG problem is solved with a cost function that includes cross terms between the state and input. They address the finite preview problem using an augmented state approach. The results in both [14] and [24] are similar to those contained in our paper. The key difference lies in the approach used to solve the problem. Specifically, our proof directly uses a version of the principle of optimality with nested information. This eliminates the need for state augmentation and hence our optimal controller is directly expressed with dynamics of the same order of the plant (not including augmented states). Moreover, our finite horizon derivation can directly accomodate non-zero initial conditions.
@@ III Problem Formulation
Consider the following discrete-time, linear time-varying (LTV) plant $P$ with the following state-space representation:
@@
where $(A,B)$ are a stabilizable pair and $A$ is nonsingular [^1] The solution to the infinite horizon problem (Section V-B) uses a related, discrete-time algebraic Riccati equation (DARE). There are technical issues in solving this DARE when $A$ is singular. Our infinite horizon results can be extended to the case where $A$ is singular using generalized solutions for the DARE. See Section 21.3 and Remaark 21.2 in [30] for details.. We will also assume a sequence of (possibly random) control policies $K\coloneq\{K_{0},K_{1},\ldots\}$ with $p$ steps of disturbance preview:
@@ IV Principle of Optimality
[…]
@@ V-A FH Stochastic LQR with Preview
Equation 30 again uses the convention that $w_{j}=0$ if $j\geq T$ [^3] Note that $\bar{v}_{t}$ depends on $\{w_{t},\ldots,w_{t+p}\}$. This means that $\bar{v}_{t+1}$ depends on $\{w_{t+1},\ldots,w_{t+p+1}\}$. However, $v_{t+1}$ defined in (26) depends on $\{w_{t+1},\ldots,w_{t+p}\}$. Hence $\bar{v}_{t+1}$ and $v_{t+1}$ differ by one term. We will show that (29) holds by induction. The base case at $t=T$ corresponds to $P_{T}=Q_{T}$, $\bar{v}_{T}=0$, and $q_{T}=0$. Next make the induction assumption that the value function at time $t+1$ has the given form:
@@
$u_{t}^{*}=\mathbb{E}\left[-K_{x,t}x_{t}-K_{w,k}w_{t}-K_{v,t}\bar{v}_{t+1}\,\middle|\,i_{t}\right].$ (33)
@@ V-C Relation to Optimal Noncausal Controller
Here the cost is deterministic (in constrast to the expected costs defined previously). The next problem states the IH design with full preview.
@@
A solution for the optimal noncausal controller is given in Theorem 11.2.1 of [29]. Related noncausal results (both FH and IH) are given in [1, 2, 3, 4, 5, 6] where the noncausal controller is used as a baseline for regret-based control design. The next result gives the IH noncausal controller in a form/notation that closely aligns with the finite preview controller in Theorem 3.
@@
Next, we show that $\|u_{t}^{nc}-u_{t}^{p}\|_{2}\to 0$ uniformly in $t$ as $p\to\infty$. Let $x_{t}^{nc}$ and $x_{t}^{p}$ denote the response of the LTI system (6) with some initial condition $x_{0}$ and disturbance $w\in\ell_{2}$ using the non-causal and $p$-step finite preview controller, respectively. The difference between the two commands is:
@@
$\begin{split}u_{t}^{nc}-u_{t}^{p}=&-K_{v}(v_{t+1}^{nc}-v_{t+1}^{p})\\ &+K_{x}\sum_{j=1}^{t}\hat{A}^{t-j}B_{u}K_{v}(v_{j}^{nc}-v_{j}^{p}).\end{split}$ (56)
@@
where $c_{1}:=\|K_{v}\|_{2\to 2}$ and $c_{2}:=\|K_{x}\|_{2\to 2}\cdot\|B_{u}K_{v}\|_{2\to 2}$ are finite constants. We can choose any $\alpha\in(\rho(\hat{A}),\,1)$ and $\exists\,T$ such that $\left\|\hat{A}^{j}\right\|\leq\alpha^{j}$ for all $j\geq T$. This follows from Gelfand’s formula (Corollary 5.6.14 in [horn2013matrix]). Therefore, the sum $\sum_{j=0}^{\infty}\|\hat{A}^{j}\|_{2\to 2}$ converges to some finite number $c_{3}$. Thus we can bound the difference in control commands as:
@@
where $K_{nc}$ is the optimal noncasual controller (51) for a specific sequence $w$. The expected per-step cost for the noncausal controller is averaged over stochastic signals $w$. We show below that it does not depend on the initial state, i.e. $J_{\infty,nc}^{*}(x_{0})=J_{\infty,nc}^{*}$. The next result is that the expected per-step cost with the optimal finite preview controller converges to the cost achieved by the optimal noncausal controller. We also give a simple expression to compute the optimal cost achieved by the noncausal controller.
@@
$\lambda_{\text{min}}(X^{\frac{1}{2}})^{2}\lambda_{\text{min}}(Y^{\frac{1}{2}})^{2}\|\hat{A}^{p+1}\|_{F}^{2}\leq J_{\infty,p}^{*}-J_{\infty,nc}^{*}\leq\|X^{\frac{1}{2}}\|_{F}^{2}\|Y^{\frac{1}{2}}\|_{F}^{2}\|\hat{A}^{p+1}\|_{F}^{2}.$
@@
$J_{\infty,p}^{*}-J_{\infty,nc}^{*}=\bigl{\|}X^{1/2}A^{p+1}Y^{1/2}\bigr{\|}_{F}^{2}.$
We obtain the upper bound in (1) using the submultiplicative property of the Frobenius norm: $\|PQ\|_{F}\leq\|P\|_{F}\,\|PQ\|_{F}$ for any two matrices $P$ and $Q$ of appropriate dimensions. We obtain the lower bound in (1) using the following property: $\|PQ\|_{F}\geq\lambda_{\text{min}}(P)\|Q\|_{F}$ where $P\succeq 0$ and $Q$ is any matrix of appropriate dimension. ∎
The spectral radius formula gives the asymptotic approximation $\lim_{k\to\infty}\|\hat{A}^{k}\|^{\frac{1}{k}}=\rho(A)$. Thus the asymptotic bounds in (1) as $p\to\infty$ are approximated by:
@@ VI Example
This section presents a simple example to illustrate the solutions to the FH and IH stochastic LQR problems. Consider a simplified LTI model for the longitudinal dynamics of a Boeing 747 linearized at one steady, level flight condition (Problem 17.6 in [35]):
@@
Figure 2 also shows the performance of the optimal non-causal controller (Theorem 4) for comparison. The cost for the optimal non-causal controller depends on the specific disturbance input. The yellow dashed line in Figure 2 corresponds to the expected (average) cost of the non-causal controller assuming white noise disturbances. This was computed using the expression given in Theorem 6. The noncausal controller has full knowledge of the disturbance and yields the lowest possible cost of $\approx 17.8$. The performance of the finite preview controller converges, as $p\to\infty$, to the performance of the optimal non-causal controller. This is expected based on the discussion in Section V-C.
Finally, Figure 3 shows the optimal IH relative error of the cost for the finite preview controller compared to the cost for the optimal noncausal controller. This is normalized by the noncausal cost and plotted on a logarithmic scale against the preview horizon $p$ (blue solid). The plot uses the same cost results as in Figure 2. The plot is linear on a log scale in agreement with the geometric convergence given in Corollary  1. The slope on the log scale should be approximately $2\log\rho(A)$. as $p\to\infty$.
@@ VII Conclusion
This paper presented solutions to the discrete-time, stochastic LQR problem with $p$ steps of disturbance preview information where $p$ is finite. Solutions were presented for both the finite horizon and infinite horizon problems. Moreover, we provided a proof for the principle of optimality that relies only on the assumption of nested information structure. This is of independent interest. Finally, we compared our solutions to several existing results for optimal control with preview information. Future work will explore using our finite-preview controller as a baseline for regret-based design.

### v5

@@ I Introduction
This paper considers control design for a discrete-time, stochastic Linear Quadratic Regulator (LQR) problem with finite preview information of the disturbance. We consider both the finite and infinite horizon problems. Our specific formulation is motivated by recent work on regret-based control [24, 25, 26, 7, 6, 5]. These regret-based formulations often use a controller with full knowledge of the disturbance as a baseline for performance. Then they show that a causal, time-varying controller, updated with online optimization, can recover this baseline performance. We show that the optimal stochastic LQR controller with finite preview can also recover this baseline performance as preview goes to infinity. Thus the performance of the noncausal controller can be recovered either via time-varying control with online optimization or using the stochastic LQR controller with an additional sensor to provide disturbance preview.
Recent work on regret-based control has also studied finite disturbance preview, e.g., in predictive or MPC-based settings [39, 14]. These works typically consider bounded disturbances and derive non-asymptotic regret bounds. In contrast, we adopt a stochastic formulation with i.i.d. disturbances. Our preview controller explicitly depends on the previewed disturbance realizations while optimizing the expected cost associated with post-preview disturbances. Rather than deriving regret bounds, we establish asymptotic convergence of the finite-preview LQR controller to the corresponding noncausal optimal controller as the preview increases.
[…]
@@
Two of the most related works are [8] and [23]. Calculus of variations is used in [8] to derive a deterministic preview control solution. The approach starts by deriving the optimal noncausal controller with full knowledge of future disturbance, similar to that given in Theorem 11.2.1 of [10]. This optimal noncausal controller is then converted to a finite-preview controller by taking the expectation for unavailable future (post-preview) information which is zero. In [23], an LQG problem is solved with a cost function that includes cross terms between the state and input. They address the finite preview problem using an augmented state approach. The results in both [8] and [23] are similar to those contained in our paper. The key difference lies in the approach used to solve the problem. Specifically, our proof directly uses a version of the principle of optimality with nested information. This eliminates the need for state augmentation and hence our optimal controller is directly expressed with dynamics of the same order of the plant (not including augmented states). Moreover, our finite horizon derivation can directly incorporate non-zero initial conditions.
@@ III Problem Formulation
Consider the discrete-time, linear time-varying (LTV) plant $P$ with the following state-space representation:
@@
where $(A,B)$ are a stabilizable pair and $A$ is nonsingular [^1] The solution to the infinite horizon problem (Section V-B) uses a related, discrete-time algebraic Riccati equation (DARE). There are technical issues in solving this DARE when $A$ is singular. Our infinite horizon results can be extended to the case where $A$ is singular using generalized solutions for the DARE. See Section 21.3 and Remark 21.2 in [40] for details.. We will also assume a sequence of (possibly random) control policies $K\coloneq\{K_{0},K_{1},\ldots\}$ with $p$ steps of disturbance preview:
@@ IV Principle of Optimality
$V_{t}^{K_{t:T-1}}(i_{t})\geq\mathbb{E}\left[\vphantom{\Big|}\mathbb{E}\left[g_{t}(x_{t},u_{t})+W_{t+1}(i_{t+1})\,\middle|\,i_{t},u_{t}=u\right]\,\middle|\,i_{t}\right].$ (17)
@@ V-A FH Stochastic LQR with Preview
Equation 30 again uses the convention that $w_{j}=0$ if $j\geq T$.  [^3] Note that $\bar{v}_{t}$ depends on $\{w_{t},\ldots,w_{t+p}\}$. This means that $\bar{v}_{t+1}$ depends on $\{w_{t+1},\ldots,w_{t+p+1}\}$. However, $v_{t+1}$ defined in (26) depends on $\{w_{t+1},\ldots,w_{t+p}\}$. Hence $\bar{v}_{t+1}$ and $v_{t+1}$ differ by one term. We will show that (29) holds by induction. The base case at $t=T$ corresponds to $P_{T}=Q_{T}$, $\bar{v}_{T}=0$, and $q_{T}=0$. Next make the induction assumption that the value function at time $t+1$ has the given form:
@@
$u_{t}^{*}=\mathbb{E}\left[-K_{x,t}x_{t}-K_{w,t}w_{t}-K_{v,t}\bar{v}_{t+1}\,\middle|\,i_{t}\right].$ (33)
@@ V-C Relation to Optimal Noncausal Controller
Here the cost is deterministic (in contrast to the expected costs defined previously). The next problem states the IH design with full preview.
@@
A solution for the optimal noncausal controller is given in Theorem 11.2.1 of [10]. Related noncausal results (both FH and IH) are given in [24, 25, 26, 7, 6] where the noncausal controller is used as a baseline for regret-based control design. The next result gives the IH noncausal controller in a form/notation that closely aligns with the finite preview controller in Theorem 3.
@@
Next, we show that $\|u_{t}^{nc}-u_{t}^{p}\|_{2}\to 0$ uniformly in $t$ as $p\to\infty$. Let $x_{t}^{nc}$ and $x_{t}^{p}$ denote the response of the LTI system (6) with some initial condition $x_{0}$ and disturbance $w\in\ell_{2}$ using the noncausal and $p$-step finite preview controller, respectively. The difference between the two commands is:
@@
$\begin{split}u_{t}^{nc}-u_{t}^{p}=-K_{v}(v_{t+1}^{nc}-v_{t+1}^{p})+K_{x}\sum_{j=1}^{t}\hat{A}^{t-j}B_{u}K_{v}(v_{j}^{nc}-v_{j}^{p}).\end{split}$ (56)
@@
where $c_{1}:=\|K_{v}\|_{2\to 2}$ and $c_{2}:=\|K_{x}\|_{2\to 2}\cdot\|B_{u}K_{v}\|_{2\to 2}$ are finite constants. We can choose any $\alpha\in(\rho(\hat{A}),\,1)$ and $\exists\,T$ such that $\left\|\hat{A}^{j}\right\|\leq\alpha^{j}$ for all $j\geq T$. This follows from Gelfand’s formula (Corollary 5.6.14 in [12]). Therefore, the sum $\sum_{j=0}^{\infty}\|\hat{A}^{j}\|_{2\to 2}$ converges to some finite number $c_{3}$. Thus we can bound the difference in control commands as:
@@
where $K_{nc}$ is the optimal noncausal controller (51) for a specific sequence $w$. The expected per-step cost for the noncausal controller is averaged over stochastic signals $w$. We show below that it does not depend on the initial state, i.e. $J_{\infty,nc}^{*}(x_{0})=J_{\infty,nc}^{*}$. The next result is that the expected per-step cost with the optimal finite preview controller converges to the cost achieved by the optimal noncausal controller. We also give a simple expression to compute the optimal cost achieved by the noncausal controller.
@@
$\lambda_{\text{min}}(X^{\frac{1}{2}})^{2}\lambda_{\text{min}}(Y^{\frac{1}{2}})^{2}\|\hat{A}^{p+1}\|_{F}^{2}\leq J_{\infty,p}^{*}-J_{\infty,nc}^{*}\leq\|X^{\frac{1}{2}}\|_{F}^{2}\|Y^{\frac{1}{2}}\|_{F}^{2}\|\hat{A}^{p+1}\|_{F}^{2}.$ (66)
@@
$J_{\infty,p}^{*}-J_{\infty,nc}^{*}=\bigl\|X^{1/2}(\hat{A})^{p+1}Y^{1/2}\bigr\|_{F}^{2}.$
We obtain the upper bound in (66) using the submultiplicative property of the Frobenius norm: $\|PQ\|_{F}\leq\|P\|_{F}\,\|Q\|_{F}$ for any two matrices $P$ and $Q$ of appropriate dimensions. We obtain the lower bound in (66) using the following property: $\|PQ\|_{F}\geq\lambda_{\text{min}}(P)\|Q\|_{F}$ where $P\succeq 0$ and $Q$ is any matrix of appropriate dimension. ∎
[…]
@@ VI Example
This section presents two numerical examples to illustrate the proposed stochastic LQR framework with disturbance preview. The first example considers a linear time-invariant (LTI) system and is included in the journal version of the paper to demonstrate the basic structure and properties of the finite-horizon and infinite-horizon solutions. The second example considers a linear time-varying (LTV) mass–spring–damper system with time-varying dynamics and cost parameters. This second example illustrates the applicability of the proposed approach to time-varying settings.
VI-A Boeing 747
Consider a simplified LTI model for the longitudinal dynamics of a Boeing 747 linearized at one steady, level flight condition (Problem 17.6 in [3]):
@@ VI-A Boeing 747
Figure 2 also shows the performance of the optimal noncausal controller (Theorem 4) for comparison. The cost for the optimal noncausal controller depends on the specific disturbance input. The yellow dashed line in Figure 2 corresponds to the expected (average) cost of the noncausal controller assuming white noise disturbances. This was computed using the expression given in Theorem 6. The noncausal controller has full knowledge of the disturbance and yields the lowest possible cost of $\approx 17.8$. The performance of the finite preview controller converges, as $p\to\infty$, to the performance of the optimal noncausal controller. This is expected based on the discussion in Section V-C.
Finally, Figure 3 shows the optimal IH relative error of the cost for the finite preview controller compared to the cost for the optimal noncausal controller. This is normalized by the noncausal cost and plotted on a logarithmic scale against the preview horizon $p$ (blue solid). The plot uses the same cost results as in Figure 2. The plot is linear on a log scale in agreement with the geometric convergence given in Corollary 1. The slope on the log scale should be approximately $2\log\rho(A)$. as $p\to\infty$.
@@
VI-B Mass-spring-damper
Consider a time-varying mass–spring–damper system with fixed damping ratio $\zeta$ and time-varying natural frequency $\omega_{n}(k)$. The system is modeled directly in discrete time as a second-order linear time-varying system
$x_{k+1}=A_{k}x_{k}+B_{k}u_{k}+E_{k}w_{k},\qquad k=0,1,\ldots,T-1,$ (69)
where $x_{k}=[\,q_{k}\;\;\dot{q}_{k}\,]^{\top}$ denotes the displacement and velocity. The time-varying system matrices are given by
$A_{k}=\begin{bmatrix}1&T_{s}\\ -T_{s}\omega_{n}(k)^{2}&1-2T_{s}\zeta\omega_{n}(k)\end{bmatrix},\qquad B_{k}=E_{k}=\begin{bmatrix}0\\ T_{s}\end{bmatrix}$.
This model corresponds to an Euler discretization of an equivalent continuous-time, mass-spring-damper system with sample time $T_{s}$. The simulation uses $T_{s}=0.02$s and $\zeta=0.7$. Moreover, the natural frequency is given by:
$w_{n}(k)=1.5+0.5\sin\left(\frac{2\pi T_{s}k}{10}\right);$ (70)
Finally, the disturbance $w_{k}$ is modeled as i.i.d. Gaussian noise.
We consider a finite-horizon stochastic LQR problem with time-varying state costs. Specifically, the input cost is fixed as $R_{k}=0.1$, while the state cost $Q_{k}$ is chosen as $\mathrm{diag}(10,1)$ in the first half of the horizon and $\mathrm{diag}(50,1)$ in the second half. This penalizes the displacement more heavily later in the horizon. We compute the optimal finite-horizon preview controller for several preview horizons $p$ and evaluate its performance using closed-loop simulations.
Figure 4 shows the average per-step cost from step 0 to step $k$ versus the step index $k$ for preview horizons $p\in\{1,5,20\}$. The performance of the optimal noncausal controller is also shown. As expected, increasing the preview horizon leads to improved performance (lower average cost). In particular, the performance with $p=20$ is already very close to that of the noncausal controller. The noncausal controller, which has access to full disturbance information, achieves the lowest cost and serves as a benchmark.
Figure 4: Comparison of average cost obtained by optimal FH controller and optimal noncausal controller in LTV system.
@@ VII Conclusion
This paper presented solutions to the discrete-time, stochastic LQR problem with $p$ steps of disturbance preview information where $p$ is finite. Solutions were presented for both the finite horizon and infinite horizon problems. Moreover, we provided a proof for the principle of optimality that relies only on the assumption of nested information structure. This is of independent interest. Finally, we compared our solutions to several existing results for optimal control with preview information. Future work will explore using our finite-preview controller as a baseline for regret-based design. Additionally, we will consider extensions to cases where the dynamic models and cost matrices are time-varying and only known with finite preview, in contrast to the full future knowledge assumed in this paper.